OpenStudy (anonymous):

Two times Owen’s age plus 4 times Charlotte’s age equals 132. Charlotte’s age is also 5 times Owen’s age. How old is Charlotte?

OpenStudy (anonymous):

if owen 's age is x charlotte 's age is y then according to above statement 2x + 4y=132 1 y = 5x 2 eq 1 becomes 2x + 4 (5x) =132 2x+20x= 132 22x= 132 x =11 putting value of x in eq 2 y =5 (11) y=55

OpenStudy (anonymous):

The floor of a rectangular tool shed has a length 10 feet greater than its width, w. Calvin will increase both dimensions of the floor by 8 feet. Which equation represents the new area, A, of the floor of the cage?